# On the combinatorics of tableaux — Classification of lattices underlying Schensted correspondences

### v1

@@ 5. Local characterization
Remark 5.20.
The poset $t$-Plait in [Fom1994a]*Exam. 2.2.10 and Fig. 7 is locally finite and cover-modular but is not graded.121212“Graded” means that every saturated chain from $a$ to $b$ has the same length, but there need not be a $\hat{0}$. In general, the rank function of a graded poset ranges over the integers, not just the nonnegative integers.
Definition 5.21.
For any $p\in J$ and any $x\in S_{p}$, we define $x^{+}$ (relative to $p$) to be the unique element of $C_{p}^{+}$ which is >x>x and we define $x^{-}$ (relative to $p$) to be the unique element of $C_{p}^{-}$ which is <x<x.
Admissible weight
Assumption 5.22.
Henceforth, we assume that $w$ is an admissible weighting, that is, the set of values $V$, the target of $w$, is $\mathbb{Z}$, and all of the values of $w$ are positive.
This assumption is not a loss of generality for our larger purpose, because if a non-admissible weighting could be used to construct an admissible weighting using th. 4.16 et al., our analysis would reveal the resulting admissible weighting directly.
@@
Assume there is $x,y\in S$, with $x\neq y$, $x^{+}=y^{+}$, and $x^{-}=y^{-}$. Since $x\neq y$, $x\not\leq y$ or $y\not\leq x$. Without loss of generality, assume $y\not\leq x$. Define $X=\{s\in S\,|\,s^{-}=x^{-}\textup{ and }s^{+}=x^{+}\textup{ and }s\leq x\}=[x]\cap[x^{-},x^{+}]$, the principal ideal generated by $x$ within the interval $x^{-},x^{+}]$. By construction, $x\in X$ and $y\not\in X$.
Because of unique-cover-modularity, given $a\in C^{-}$, $a\leq z$ for some $z\in X$ iff $a=x^{-}$ iff $a\leq x$. So $a\not\leq X$ iff $a\not\leq x$ Because $y\in S\setminus X$ and $y^{+}=x^{+}$, given $b\in C^{+}$, $b\geq z$ for some $z\in S$ iff $b\geq z$ for some $z\in S\setminus X$. So $b\not\geq S$ iff $b\not\geq S\setminus X$.
Applying (14) to $S=\emptyset\sqcup S$ gives
$\sum_{a\in C^{-},\ a\not\leq\emptyset}w(a)+\sum_{b\in C^{+},\ b\not\geq S}w(b)$ $=2w(p)$
$\sum_{a\in C^{-}}w(a)+\sum_{b\in C^{+},\ b\not\geq S\setminus X}w(b)$ $=2w(p)$
Applying (14) to $S=X\sqcup(S\setminus X)$ gives
$\sum_{a\in C^{-},\ a\not\leq X}w(a)+\sum_{b\in C^{+},\ b\not\geq S\setminus X}w(b)$ $=2w(p)$
$\sum_{a\in C^{-},\ a\not\leq x}w(a)+\sum_{b\in C^{+},\ b\not\geq S\setminus X}w(b)$ $=2w(p)$
Taking the difference of these gives $w(x^{-})=0$. By the assumption that $w$ is admissible, this is a contradiction. ∎
@@
Assume there exists $x<y$ in $S_{p}$. By unique-cover-modularity, $x^{+}=y^{+}$ and $x^{-}=y^{-}$. By lem. 5.23, $x=y$, which is a contradiction. Thus $S_{p}$ is an antichain. ∎
Thus, for any $p$, any subset of $S_{p}$ is both an ideal and a filter, which allows (14) and (17) to be used quite broadly.
Consider how unique-cover-modularity simplifies (17): For any given $m\in S$, the $a$ on the left-hand side of (17) can only be $m^{-}$. Similarly, the $b$ on the right-hand side can only be $m^{+}$. Thus (17) is equivalent to:
$\begin{rcases}\textup{if }m^{-}\not\leq L\textup{ }&w(m^{-})\\ \textup{otherwise }&0\end{rcases}$ $=\begin{cases}w(m^{+})&\textup{ if }m^{+}\not\geq U\\ 0&\textup{ otherwise}\end{cases}$ (20)
If we set $L=\emptyset$ and $U=S\setminus\{m\}$,
$w(m^{-})$ $=\begin{cases}w(m^{+})&\textup{ if }m^{+}\not\geq U\\ 0&\textup{ otherwise}\end{cases}$ (21)
Given the assumption of unique-cover-modularity, $m^{+}\geq z$ for some $z\in U$ iff $m^{+}=z^{+}$ for some $z\in U$. So $m^{+}\not\geq U$ iff there is no $z\in U$ with $m^{+}=z+$. Given the assumption of admissibility, $w(m^{-})>0$, which is contradicted by (21) unless there is no $z\in S$ with $z^{+}=m^{+}$ and $z\neq m$.
Dually, if we set $U=\emptyset$ and $L=S\setminus\{m\}$, we get a contradiction unless there is no $z\in S$ with $z^{-}=m^{-}$ and $z\neq m$.
Thus for every $m\neq n$ in $S$, $m^{-}\neq n^{-}$ and $m^{+}\neq n^{+}$; the set of pairs $\{(m^{-},m^{+})\,|\,m\in S\}$ is a bijection between subsets of $C^{-}$ and $C^{+}$. In addition, (20) reduces to
$w(m^{-})$ $=w(m^{+})\textup{\qquad for all $m\in S_{p}$, $p\in J$}$ (22)
Note that $x^{+}$ and $x^{-}$ in (22) implicitly depend on $p$.
@@
Given a $p\in J$, (17) holds for every partition of $S_{p}$ into three disjoint sets $S_{p}=L\sqcup\{m\}\sqcup U$ iff the following conditions are true for $p$:
@@
    the set of pairs $\{(m^{-},m^{+})\,|\,m\in S_{p}\}$ is a bijection between subsets of $C^{-}$ and $C^{+}$ and
@@
Regarding $\Leftarrow$: By the above derivation, for any $S_{p}=L\sqcup\{m\}\sqcup U$, (17) is the same as (20). By hypothesis (1), $m^{-}\leq z$ for some $z\in L$ iff $m^{-}=z^{-}$ iff $m=z$. But by construction, $m\not\in L$. Thus $m^{-}\not\leq L$ is always true. Dually, $m^{+}\not\leq U$ is always true. Thus (20) is the same as
$w(m^{-})$ $=w(m^{+})$ (23)
which is assumed as hypothesis (2). ∎
@@
  1. (assump. 5.1) $L$ is a Fomin lattice,
@@
  3. (assump. 5.1) $L$ is distributive, and
  4. (assump. 5.19) $J$ is unique-cover-modular,
a weighting $w(\bullet\lessdot\bullet)$ of coverings in $L$ into $V=\mathbb{Z}$ satisfies the differential condition (7) and is admissible (assump. 5.22) iff the corresponding $w(\bullet)$ from $J$ into $V$ satisfies
  1. (assump. 5.22) $w(p)>0$ for every $p\in J$,
  2. (eq. 11) $r=\sum_{z\,|\,z\textup{ minimal in }J}w(z)$,
  3. (eq. 15) $\sum_{a\,|\,a\lessdot p}w(a)+\sum_{b\,|\,b\gtrdot p,\ b\textup{ upward orphan}}w(b)=2w(p)$ for all $p\in J$,
  4. (lem. 5.25(1)) the set of pairs $\{(m^{-},m^{+})\,|\,m\in S_{p}\}$ is a bijection between subsets of $C^{-}_{p}$ and $C^{+}_{p}$ for all $p\in J$,
  5. (lem. 5.25(2)) $w(m^{-})=w(m^{+})$ for every $m\in S_{p}$, $p\in J$.
@@
This is proved by chaining together th. 5.2, lem. 5.7, lem. 5.13, lem. 5.14, lem. 5.18, and lem. 5.25. ∎
@@
No element of $J$ is covered by three or more distinct elements.
@@
Figure 4. Construction for the proof of lem. 5.27
Assume that we have $a\in J$ covered by distinct $p_{1}$, $q_{1}$ and $r_{1}$. By unique-cover-modularity, there is a $pq_{2}$ covering $p_{1}$ and $q_{1}$, a $qr_{2}$ covering $q_{1}$ and $r_{1}$, and an $rp_{2}$ covering $r_{1}$ and $p_{1}$. By lem. 5.23 (applied to $p_{1}$, $q_{1}$, and $r_{1}$), $pq_{2}$, $qr_{2}$, and $rp_{2}$ are distinct. By (22) (applied to $p_{1}$, $q_{1}$, and $r_{1}$), $w(a)=w(pq_{2})=w(qr_{2})=w(rp_{2})$.
By unique-cover-modularity, there is a $p_{3}$ covering $rp_{2}$ and $pq_{2}$, a $q_{3}$ covering $pq_{2}$ and $qr_{2}$, and an $r_{3}$ covering $qr_{3}$ and $rp_{3}$. Suppose two of $p_{3}$, $q_{3}$, and $r_{3}$ are the same, say $p_{3}$ is the same as $q_{3}$. Define $b=p_{3}=q_{3}$, so $b$ covers $pq_{2}$, $qr_{2}$, and $rp_{2}$. Then by (22) (applied to $pq_{2}$, $qr_{2}$, and $rp_{2}$), $w(b)=w(p_{1})=w(q_{1})=w(r_{1})$. Since $w$ is admissible, (15) implies $2w(pq_{2})\geq w(p_{1})+w(q_{1})=2w(b)$. So $w(pq_{2})\geq w(b)$ and similarly $w(qr_{2})\geq w(b)$ and $w(rp_{2})\geq w(b)$. Applying (15) to $b$ and the above gives $2w(b)\geq w(pq_{2})+w(qr_{2})+w(rp_{2})\geq 3w(b)$. Since $w(b)>0$, this is a contradiction. Thus $p_{3}$, $q_{3}$, and $r_{3}$ are distinct.
By (22), $w(p_{3})=w(p_{1})$, $w(q_{3})=w(q_{1})$, and $w(r_{3})=w(r_{1})$. Applying (15) and admissibility,
@@
Since the inequalities (24) and (25) are tight, (15) and the assumption that $w$ is admissible requires that $pq_{2}$, $qr_{2}$, $rp_{2}$, $p_{3}$, $q_{3}$, and $r_{3}$ cover no elements other than the ones we mentioned above (and shown in fig. 4) and have no upward orphans. Applying (15) to $pq_{2}$, $qr_{2}$, and $rp_{2}$ thus gives
@@
By (16) and that $w$ is admissible, $2w(a)\geq w(p_{1})+w(q_{1})+w(r_{1})=3w(a)$. Since $w(a)>0$, this is a contradiction.
Thus, there is no $a\in J$ that is covered by three distinct elements. ∎
@@
No element of $J$ covers three or more distinct elements.
@@
This is proved dually to lem. 5.27, by interchanging the use of (15) and (16) and replacing “upward orphan” with “downward orphan”. ∎
@@
Lem. 5.27 and 5.28 are the deeper explanation of the long-known fact[Fom1994a]*sec. 2.2 that ${\operatorname{Ideal}_{\textup{f}}}(\mathbb{N}^{k})$ for $k>2$ has no admissible weighting as a differential lattice. Similarly, they show why $\mathbb{SS}_{t}={\operatorname{Ideal}_{\textup{f}}}(\mathbb{CP}_{2t})={\operatorname{Ideal}_{\textup{f}}}(t\textit{-SkewStrip})$ is a Fomin lattice (para. 2(9) but ${\operatorname{Ideal}_{\textup{f}}}(\mathbb{CP}_{kt})$ with $k>2$ has no admissible weightings. (Compare $\mathbb{CP}_{32}=\mathcal{L}_{3,2}$ in [El2025a]*Fig. 2 with $\mathbb{CP}_{33}=\mathcal{L}_{3,3}$ in [El2025a]*Fig. 3.)
The equations implied by lem. 5.26 for any particular $p\in J$ are characterized by the structure of its neighborhood $N_{p}$ as a poset.
@@
$N_{p}$ is characterized up to isomorphism as a poset by the numbers $|C^{-}|$, $|C^{+}|$ and $|S|$: All elements of $C^{-}$ are covered by $p$, all elements of $C^{+}$ cover $p$, and $S$ forms a matching between subsets of $C^{-}$ and $C^{+}$, with each element of $S$ covering a distinct element of $C^{-}$ and being covered by a distinct element of $C^{+}$. In addition, $|C^{-}|\leq 2$, $|C^{+}|\leq 2$, and $|N|=|C^{-}|+|C^{+}|+|S|+1$.
@@
This follows from lem. 5.24, 5.26, 5.27, and 5.28. ∎
For each possible set of values of $|C^{-}|$, $|C^{+}|$, $|S|$, and $|N|$, fig. 5 shows the resulting structure of $N$, and the consequent constraints per lem. 5.26.
$|C^{+}|$ $|C^{-}|$ $|S|$ $|N|$ Neighborhood Constraints

### v2

@@ 5. Local characterization
Henceforth, we assume that $w$ is an positive weighting, that is, the set of values $V$, the target of $w$, is $\mathbb{Z}$, and all of the values of $w$ are positive.
This assumption is not a loss of generality for our larger purpose, because if a non-positive weighting could be used to construct a positive weighting using th. 4.16 et seq., our analysis would reveal the resulting positive weighting directly.
@@
Assume there is $x,y\in S_{p}$, with $x\neq y$, $x^{+}_{p}=y^{+}_{p}$, and $x^{-}_{p}=y^{-}_{p}$. Since $x\neq y$, either $x\not\leq y$ or $y\not\leq x$. Without loss of generality, assume $y\not\leq x$. Define $X=\{s\in S_{p}\,|\,s^{-}_{p}=x^{-}_{p}\textup{ and }s^{+}_{p}=x^{+}_{p}\textup{ and }s\leq x\}=[x]\cap[x^{-}_{p},x^{+}_{p}$], the principal ideal generated by $x$ within the interval $x^{-}_{p},x^{+}_{p}$]. By construction, $x\in X$ and $y\not\in X$.
Given $a\in C_{p}^{-}$, $a\leq X$ iff $a\leq x$, so $a\not\leq X$ iff $a\not\leq x$. Given $b\in C_{p}^{+}$, if $b\geq$ some $z\in X$, $z^{+}_{p}=b$. But since $z\leq x$, lem. 5.15 shows $z^{+}_{p}=x^{+}_{p}$, so $b=z^{+}_{p}=x^{+}_{p}=y^{+}_{p}$ and $b\geq y\not\in X$. Thus $b\geq S_{p}$ iff $b\geq S_{p}\setminus X$ and $b\not\geq S_{p}$ iff $b\not\geq S_{p}\setminus X$.
Applying (14) to $S_{p}=\emptyset\sqcup S_{p}$ gives
$\sum_{a\in C_{p}^{-},\ a\not\leq\emptyset}w(a)+\sum_{b\in C_{p}^{+},\ b\not\geq S_{p}}w(b)$ $=2w(p)$
which is equivalent to
$\sum_{a\in C_{p}^{-}}w(a)+\sum_{b\in C_{p}^{+},\ b\not\geq S_{p}\setminus X}w(b)$ $=2w(p)$ (22)
Applying (14) to $S_{p}=X\sqcup(S_{p}\setminus X)$ gives
$\sum_{a\in C_{p}^{-},\ a\not\leq X}w(a)+\sum_{b\in C_{p}^{+},\ b\not\geq S_{p}\setminus X}w(b)$ $=2w(p)$
which is equivalent to
$\sum_{a\in C_{p}^{-},\ a\not\leq x}w(a)+\sum_{b\in C_{p}^{+},\ b\not\geq S_{p}\setminus X}w(b)$ $=2w(p)$ (23)
Subtracting (23) from (22) gives $w(x^{-}_{p})=0$. By the assumption that $w$ is positive, this is a contradiction. Thus $x=y$. ∎
@@
Assume there exists $x<y$ in $S_{p}$. By lem. 5.15, $x^{+}_{p}=y^{+}_{p}$ and $x^{-}_{p}=y^{-}_{p}$. By lem. 5.25, $x=y$, which is a contradiction. Thus $S_{p}$ is an antichain. ∎
Thus, for any $p$, any subset of $S_{p}$ is both an ideal and a filter, which allows (14) and (19) to be applied quite broadly. Consider how unique-cover-modularity simplifies (19): For any given $m\in S_{p}$, any $a$ on the left-hand side of (19) can only be $m^{-}_{p}$. Similarly, any $b$ on the right-hand side can only be $m^{+}_{p}$. Thus (19) is equivalent to:
$\begin{rcases}\textup{if }m^{-}_{p}\not\leq L\textup{ }&w(m^{-}_{p})\\ \textup{otherwise }&0\end{rcases}$ $=\begin{cases}w(m^{+}_{p})&\textup{ if }m^{+}_{p}\not\geq U\\ 0&\textup{ otherwise}\end{cases}$ (24)
If we set $L=\emptyset$ and $U=S_{p}\setminus\{m\}$,
$w(m^{-}_{p})$ $=\begin{cases}w(m^{+}_{p})&\textup{ if }m^{+}_{p}\not\geq U\\ 0&\textup{ otherwise}\end{cases}$ (25)
Given the assumption of unique-cover-modularity, $m^{+}_{p}\geq z$ for some $z\in U$ iff $m^{+}_{p}=z^{+}_{p}$ for some $z\in U$. So $m^{+}_{p}\not\geq U$ iff there is no $z\in U$ with $m^{+}_{p}=z^{+}_{p}$. Given the assumption of positivity, $w(m^{-}_{p})>0$, which is contradicted by (25) unless there is no $z\in S_{p}$ with $z^{+}_{p}=m^{+}_{p}$ and $z\neq m$.
Dually, if we set $L=S_{p}\setminus\{m\}$ and $U=\emptyset$, we get a contradiction unless there is no $z\in S_{p}$ with $z^{-}_{p}=m^{-}_{p}$ and $z\neq m$.
Thus for every $m\neq n$ in $S_{p}$, $m^{-}_{p}\neq n^{-}_{p}$ and $m^{+}_{p}\neq n^{+}_{p}$, and so the set of pairs $\{(m^{-}_{p},m^{+}_{p})\,|\,m\in S_{p}\}$ is a bijection between subsets of $C_{p}^{-}$ and $C_{p}^{+}$. In addition, (24) reduces to
$w(m^{-}_{p})$ $=w(m^{+}_{p})\textup{\qquad for all $p\in P$ and $m\in S_{p}$}$ (26)
@@
Given $p\in P$, (19) holds for every partition of $S_{p}$ into three disjoint sets $S_{p}=L\sqcup\{m\}\sqcup U$ with $L$ an ideal of $S_{P}$ and $U$ a filter of $S_{p}$ iff the following conditions are true:
@@
    the set of pairs $\{(m^{-}_{p},m^{+}_{p})\,|\,m\in S_{p}\}$ is a bijection between subsets of $C_{p}^{-}$ and $C_{p}^{+}$ and
@@
[…]
@@
  1. (assump. 5.1) $L$ satisfies the criteria of a Fomin lattice other than $(A)$, the weighted-differential property,
@@
  3. (assump. 5.1) $L$ is distributive,
  4. (assump. 5.13) $P$ is unique-cover-modular, and
  5. (assump. 5.24) $w(\bullet)$ from $P$ into $V=\mathbb{Z}$ has only positive values,
then $L$ satisfies (9) (and thus is a Fomin lattice) iff
  1. (eq. 11) $r=\sum_{z\,|\,z\textup{ minimal in }P}w(z)$,
  2. (eq. 15) $\sum_{a\,|\,a\lessdot p}w(a)+\sum_{b\,|\,b\gtrdot p,\ b\textup{ upward orphan}}w(b)=2w(p)$ for all $p\in P$,
  3. (lem. 5.27(1)) the set of pairs $\{(m^{-}_{p},m^{+}_{p})\,|\,m\in S_{p}\}$ is a bijection between subsets of $C_{p}^{-}$ and $C_{p}^{+}$ for all $p\in P$, and
[…]
@@
This is proved by chaining together th. 5.2, lem. 5.7, lem. 5.17, lem. 5.18, lem. 5.22, and lem. 5.27. ∎
Similarly to lem. 5.22 and 5.23, we can replace (15) with (16):
Lemma 5.29.
Given
  1. (assump. 5.1) $L$ satisfies the criteria of a Fomin lattice other than $(A)$, the weighted-differential property,
  2. (assump. 5.1) $L$ has a $\hat{0}$,
  3. (assump. 5.1) $L$ is distributive,
  4. (assump. 5.13) $P$ is unique-cover-modular, and
  5. (assump. 5.24) $w(\bullet)$ from $P$ into $V=\mathbb{Z}$ has only positive values,
then $L$ satisfies (9) (and thus is a Fomin lattice) iff
  1. (eq. 11) $r=\sum_{z\,|\,z\textup{ minimal in }P}w(z)$,
  2. (eq. 16) $\sum_{a\,|\,a\lessdot p,\ a\textup{ downward orphan}}w(a)+\sum_{b\,|\,b\gtrdot p}w(b)=2w(p)$ for all $p\in P$,
  3. (lem. 5.27(1)) the set of pairs $\{(m^{-}_{p},m^{+}_{p})\,|\,m\in S_{p}\}$ is a bijection between subsets of $C_{p}^{-}$ and $C_{p}^{+}$ for all $p\in P$, and
[…]
Proof.
This is proved by chaining together th. 5.2, lem. 5.7, lem. 5.17, lem. 5.18, lem. 5.23, and lem. 5.27. ∎
@@
No element of $P$ is covered by three or more distinct elements.
@@
Figure 8. Construction for the proof of lem. 5.30
Assume that we have $a\in P$ covered by distinct $p_{1}$, $q_{1}$ and $r_{1}$. By unique-cover-modularity, there is a $pq_{2}$ covering $p_{1}$ and $q_{1}$, a $qr_{2}$ covering $q_{1}$ and $r_{1}$, and an $rp_{2}$ covering $r_{1}$ and $p_{1}$. By lem. 5.25 (applied to $p_{1}$, $q_{1}$, and $r_{1}$), $pq_{2}$, $qr_{2}$, and $rp_{2}$ are distinct. By (26) (applied to $p_{1}$, $q_{1}$, and $r_{1}$), $w(a)=w(pq_{2})=w(qr_{2})=w(rp_{2})$.
By unique-cover-modularity, there is a $p_{3}$ covering $rp_{2}$ and $pq_{2}$, a $q_{3}$ covering $pq_{2}$ and $qr_{2}$, and an $r_{3}$ covering $qr_{3}$ and $rp_{3}$. Suppose two of $p_{3}$, $q_{3}$, and $r_{3}$ are the same, say $p_{3}$ is the same as $q_{3}$. Define $b=p_{3}=q_{3}$, so $b$ covers $pq_{2}$, $qr_{2}$, and $rp_{2}$. Then by (26) (applied to $pq_{2}$, $qr_{2}$, and $rp_{2}$), $w(b)=w(p_{1})=w(q_{1})=w(r_{1})$. Since $w$ is positive, (15) implies $2w(pq_{2})\geq w(p_{1})+w(q_{1})=2w(b)$. So $w(pq_{2})\geq w(b)$ and similarly $w(qr_{2})\geq w(b)$ and $w(rp_{2})\geq w(b)$. Applying (15) to $b$ and the above gives $2w(b)\geq w(pq_{2})+w(qr_{2})+w(rp_{2})\geq 3w(b)$. Since $w(b)>0$, this is a contradiction. Thus $p_{3}$, $q_{3}$, and $r_{3}$ are distinct.
By (26) (applied to $pq_{2}$, $qr_{2}$, and $rp_{2}$), $w(p_{3})=w(p_{1})$, $w(q_{3})=w(q_{1})$, and $w(r_{3})=w(r_{1})$. Applying (15) and positivity,
@@
and the inequalities (27) and (28) are equalities. Equation (15) and the assumption that $w$ is positive requires that $pq_{2}$, $qr_{2}$, $rp_{2}$, $p_{3}$, $q_{3}$, and $r_{3}$ cover no elements other than the ones mentioned above (and shown in fig. 8) and have no upward orphans. Applying (15) to $pq_{2}$, $qr_{2}$, and $rp_{2}$ thus gives
@@
By (16) and that $w$ is positive, $2w(a)\geq w(p_{1})+w(q_{1})+w(r_{1})=3w(a)$. Since $w(a)>0$, this is a contradiction.
Thus, there is no $a\in P$ that is covered by three distinct elements. ∎
@@
No element of $P$ covers three or more distinct elements.
@@
This is proved dually to lem. 5.30, by interchanging the use of (15) and (16) and replacing “upward orphan” with “downward orphan”. ∎
@@
Lem. 5.30 and 5.31 are the deeper explanation of the long-known fact[Fom1994a]*sec. 2.2 that ${\operatorname{Ideal}_{\textup{f}}}(\mathbb{N}^{k})$ for $k>2$ has no positive weighting as a differential lattice. Similarly, they show why $\mathbb{SS}_{t}={\operatorname{Ideal}_{\textup{f}}}(\mathbb{CP}_{2t})={\operatorname{Ideal}_{\textup{f}}}(t\textit{-SkewStrip})$ is a Fomin lattice (para. 2 item (10)) but ${\operatorname{Ideal}_{\textup{f}}}(\mathbb{CP}_{kt})$ with $k>2$ has no positive weightings. (Compare $\mathbb{CP}_{32}=\mathcal{L}_{3,2}$ in [El2025a]*Fig. 2 with $\mathbb{CP}_{33}=\mathcal{L}_{3,3}$ in [El2025a]*Fig. 3.)
The equations implied by lem. 5.28 for any particular $p\in P$ are characterized by the structure of its neighborhood $N_{p}$ as a poset.
@@
$N_{p}$ is characterized up to poset isomorphism fixing $p$ by the numbers $|C_{p}^{-}|$, $|C_{p}^{+}|$ and $|S_{p}|$: All elements of $C_{p}^{-}$ are covered by $p$, all elements of $C_{p}^{+}$ cover $p$, and $S_{p}$ forms a matching between subsets of $C_{p}^{-}$ and $C_{p}^{+}$, with each element of $S_{p}$ covering a distinct element of $C_{p}^{-}$ and being covered by a distinct element of $C_{p}^{+}$. In addition, $|C_{p}^{-}|\leq 2$, $|C_{p}^{+}|\leq 2$, $|S_{p}|\leq|C_{p}^{+}|$, $|S_{p}|\leq|C_{p}^{-}|$, and $|N_{p}|=|C_{p}^{-}|+|C_{p}^{+}|+|S_{p}|+1$.
@@
This follows from lem. 5.26, 5.27, 5.30, and 5.31. ∎
For each possible set of values of $|C_{p}^{-}|$, $|C_{p}^{+}|$, $|S_{p}|$, and $|N_{p}|$, fig. 5 shows the resulting structure of $N_{p}$, and the consequent constraints per lem. 5.28.
$|C_{p}^{+}|$ $|C_{p}^{-}|$ $|S_{p}|$ $|N_{p}|$ Neighborhood Constraints